OpenStudy (anonymous):

how do you solve this √(3x-1) =6

OpenStudy (anonymous):

\[\sqrt{(3x-1)}=6\] square both sides \[(\sqrt{(3x-1)})^2 = (6)^2\] \[3x-1 = 36\] add one to both sides \[3x = 37\] divide both sides by 3 \[x = 37/3\]

OpenStudy (anonymous):

if that helps

OpenStudy (anonymous):

thank u I was trying to help my son with his homework he has been at the table for hours I just came home from work an hour ago and he was so frustrated truly i thank u